# Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback

### v1

@@ Abstract
We consider offline reinforcement learning (RL) with preference feedback in which the implicit reward is a linear function of an unknown parameter. Given an offline dataset, our objective consists in ascertaining the optimal action for each state, with the ultimate goal of minimizing the simple regret. We propose an algorithm, RL with Locally Optimal Weights or RL-LOW, which yields a simple regret of $\exp(-\Omega(n/H))$ where $n$ is the number of data samples and $H$ denotes an instance-dependent hardness quantity that depends explicitly on the suboptimality gap of each action. Furthermore, we derive a first-of-its-kind instance-dependent lower bound in offline RL with preference feedback. Interestingly, we observe that the lower and upper bounds on the simple regret match order-wise in the exponent, demonstrating order-wise optimality of RL-LOW. In view of privacy considerations in practical applications, we also extend RL-LOW to the setting of $(\varepsilon,\delta)$-differential privacy and show, somewhat surprisingly, that the hardness parameter $H$ is unchanged in the asymptotic regime as $n$ tends to infinity; this underscores the inherent efficiency of RL-LOW in terms of preserving the privacy of the observed rewards. Given our focus on establishing instance-dependent bounds, our work stands in stark contrast to previous works that focus on establishing worst-case regrets for offline RL with preference feedback.
@@ 1 Introduction
Reinforcement Learning (RL) (Sutton and Barto,, 2018) has been widely recognized for its capacity to facilitate agents in learning a sequence of optimal actions through iterative interactions with their environments. However, RL encounters significant hurdles in environments that are characterized by uncertainty or lacking explicit reward signals. To address these shortcomings, the concept of RL with human feedback (or RLHF) has emerged as a prominent paradigm. Preference-based RL (PbRL) (Christiano et al.,, 2017; Chen et al.,, 2022; Ibarz et al.,, 2018; Palan et al.,, 2019) has stood out as one of the most widely used frameworks for RLHF. In this regard, preference-based RL has achieved remarkable performances in practical applications, with particular importance lying in its ability to align large language models (LLMs) with human intent, thereby mitigating the output of toxic and dishonest information (Ouyang et al.,, 2022; Ziegler et al.,, 2019; Glaese et al.,, 2022; Bai et al.,, 2022; Liu et al.,, 2023), and improving the quality of applying to the specific tasks (Stiennon et al.,, 2020; Wu et al.,, 2021; Nakano et al.,, 2021).
@@
In this work, we tackle the problem of offline RL with preference feedback, wherein the learning mechanism operates solely on pre-existing (or offline) data without dynamically engaging with the environment. Given the high cost associated with human interaction, offline RL has assumed particular importance in the context of incorporating human feedback. The significance of this offline framework has been justified by many previous prominent works (Shin et al.,, 2023; Ouyang et al.,, 2022; Zhu et al.,, 2023; Kim et al.,, 2023). For instance, within the learning process of InstructGPT (Ouyang et al.,, 2022), a pivotal procedure involves the training of a reward model utilizing pre-trained LLM feature vectors, coupled with the utilization of pre-collected human preference feedback as the training dataset. Conceptually, this procedure can be construed as treating the current prompt context as a state within a certain Markov Decision Process (MDP), while the responses generated by the LLM serve as actions within this process. Empirical findings presented by Ouyang et al., (2022) demonstrate the efficacy of this offline framework in effectively aligning human intent with the outputs of LLMs.
@@
However, the literature concerning theoretical analyses within the domain of offline PbRL remains rather scant. Previous theoretical analyses (Zhu et al.,, 2023; Zhan et al.,, 2024) of offline PbRL predominantly focused on the worst-case (or minimax) regret, often resulting in the derivation of regret upper bounds for their algorithms of the form $\tilde{O}(n^{-1/2})$, where $n$ is the size of the offline dataset. In this work, we adopt a different approach that is centered on instance-dependent guarantees. In other words, we wish to derive performance guarantees that are functions of the specific problem instance, thus elucidating the role of fundamental hardness parameters. This yields complementary insights to the existing worst-case analyses. To this end, we design and analyze RL-LOW, a preference-based RL algorithm. Our analysis of the performance RL-LOW unveils an instance-dependent simple regret bound of $\exp(-\Omega(n/H))$, where $H$ is a hardness parameter. This reveals that the simple regret decays exponentially fast in the size of the dataset $n$ and the exponential rate of convergence has also been identified. Complementarily, by proving an instance-dependent lower bound, we show that any algorithm will suffer from a simple regret of at least $\exp(-O(n/H))$. Thus, the dependence of the problem on $H$ is fundamental and cannot be improved upon, thereby demonstrating the efficacy of RL-LOW and the tightness of our analyses.
@@ 1.1 Related works
Preference-based RL: From the empirical viewpoint, Christiano et al., (2017) initially demonstrated that RL systems can effectively address complex tasks like Atari games and simulated robot locomotion by learning from human preferences between trajectory segments. Later, numerous researchers started to employ human preference feedback to enhance the performance of LLMs, e.g., aligning the LLMs’ behavior with human intent (Ouyang et al.,, 2022; Ziegler et al.,, 2019; Glaese et al.,, 2022; Bai et al.,, 2022; Liu et al.,, 2023), and enhancing the efficacy of application to specific tasks (Stiennon et al.,, 2020; Wu et al.,, 2021; Nakano et al.,, 2021).
@@
From the theoretical perspective, the existing literature remains sparse in offline RL with preference feedback. Zhu et al., (2023) elucidated the failure of the maximum likelihood estimation (MLE) procedure in some scenarios. Motivated by this, they theoretically prove the (near) minimax optimality of the Pessimistic MLE approach with a high probability guarantee. In addition,  Zhan et al., (2024) introduced a novel paradigm for general reward functions, and they introduce $\varepsilon$-bracket approximations for reward models, accompanied by a rigorous theoretical analysis delineating sample complexity in terms of approximation error $\varepsilon$ and the high-probability parameter $\delta$.
@@
We observe that the above theoretical investigations, while invaluable, are not instance-dependent. Typically, the above minimax or worst-case guarantees yield upper bounds in the form of $\tilde{O}(n^{-1/2})$ and do not depend on any problem-specific factors (such as suboptimality gaps). Our research stands out as a pioneering attempt in offering an instance-dependent examination for offline RL with preference feedback, thereby bridging a critical gap in the existing literature.
@@
Label-Differential Privacy: In our study, we also consider the notion of label privacy, acknowledging that the labels in our offline dataset originate from users, thus highlighting the imperative to protect user privacy. Chaudhuri and Hsu, (2011) were among the pioneers in exploring the concept of label privacy within the context of supervised learning for binary classification. Their foundational work posits that the sensitive information primarily resides in the labels, while considering the unlabeled attributes as non-sensitive. Later, the concept of label privacy has been investigated across various machine learning paradigms, including but not limited to PAC learning (Beimel et al.,, 2013) and deep learning frameworks (Ghazi et al.,, 2021). This broadened examination underscores the significance and relevance of label privacy considerations across diverse areas of machine learning research and applications.
@@
More recently, Chowdhury et al., (2024) investigated the use of label differential privacy to protect the privacy of human labelers in the process of estimating rewards from preference-based feedback. Chowdhury et al., (2024) derive an upper bound for their proposed algorithm on the estimation error. They show that it also decays as $O(n^{-1/2})$ and the implied constant here depends on $(\varepsilon,\delta)$, the parameters that define differential privacy. This bound only applies in the scenario of estimating the reward value and is not applicable if we want to understand how it depends on the simple regret of a specific instance. In our work, we consider the effect of $(\varepsilon,\delta)$-DP on the simple regret.
@@ 1.2 Our contributions
    We adapt RL-LOW to the $(\varepsilon,\delta)$-differential privacy paradigm by combining the Gaussian mechanism with the aforementioned locally optimal weights. Our analysis demonstrates that, for large datasets, this combination enables our algorithm to achieve differential privacy without weakening the bound on the simple regret, underscoring the superiority of the design and analysis of RL-LOW.
@@
    As a by-product of our analyses, we show that RL-LOW achieves a worst-case bound of the form $O(n^{-1/2})$. If we translate the high-probability upper bound in Zhu et al. Zhu et al., (2023) to the same worst-case setting, we obtain a bound of the form $O(\sqrt{n^{-1}\log n})$. Thus, our work provides a noticeable (albeit small) improvement over the state-of-the-art theoretical result in Zhu et al., (2023).
@@ 2 Preliminaries and Problem Setup
Let $\mathcal{S}=\{1,\ldots,S\}$ denote the state space, and $\mathcal{A}=\{1,\ldots,A\}$ denote the action set. We assume that there is an unknown non-degenerate distribution $\rho=(\rho_{1},\ldots,\rho_{S})$ over the states, i.e., $\rho_{k}>0$ for all $k\in\mathcal{S}$. The $i$-th action of state $k$ is associated with the feature vector $\phi(k,i)\in\mathbb{R}^{d}$, and its associated (unknown) reward is
@@
where $\theta\in\mathbb{R}^{d}$ is an unknown parameter vector. The collection of all feature vectors is denoted as $\phi=\{\phi(k,i)\}_{k\in\mathcal{S},i\in\mathcal{A}}$. For all $k\in\mathcal{S}$, we denote the suboptimaliy gap of action $i\in\mathcal{A}$ as $\Delta_{k,i}=\max_{j\in\mathcal{A}}r_{k,j}-r_{k,i}$. Let $(a^{(0)},a^{(1)})\in\mathcal{A}^{2}$ be a pair of comparisons and let $s\in\mathcal{S}$ be a state. Then, we define a stochastic label $\sigma$, following the Bradley–Terry–Luce (BTL) model as
@@
Given this model, we assume throughout that we have access to an offline dataset, which we denote as $\mathcal{D}=\{({s}_{i},{a}^{(0)}_{i},{a}^{(1)}_{i},\sigma_{i})\}_{i=1}^{n}$. Note that this dataset consists of $n$ tupies of states, pairs of actions for comparison, and stochastic labels. Without loss of generality, we assume that the comparisons are arranged such that $a^{(0)}_{i}<a^{(1)}_{i}$ for all $i=1,\ldots,n$, and $a^{(0)}_{i}<a^{(0)}_{j}$ (or $a^{(1)}_{i}\leq a^{(1)}_{j}$ if $a^{(0)}_{i}=a^{(0)}_{j}$) for all $i<j$. For simplicity, we assume that the feature vectors satisfy $\phi(k,i)\neq\phi(k,j)$ for all states $k\in\mathcal{S}$ and all actions $i\neq j$. In addition, we assume that for each state $k$, the best action $i_{k}^{*}=\arg\max_{j\in\mathcal{A}}r_{k,j}$ is unique. Broadly speaking, our objective is to use the offline dataset $\mathcal{D}$ to estimate the best action $i_{k}^{*}$ for each state $k\in\mathcal{S}$. Following  Zhu et al., (2023), we aim to design a (possibly randomised) policy $\pi$ that uses the dataset $\mathcal{D}$ to output a set of actions $\{\hat{i}_{k}\}_{k\in\mathcal{S}}$ that minimizes the simple regret [^1] The term “simple regret” is referred to as “performance gap” in some existing works (e.g., Zhu et al., (2023)). (averaged over all $S$ states), defined as
$R_{n}=\sum_{k\in\mathcal{S}}\rho_{k}\big{(}r_{k,i^{*}_{k}}-r_{k,\hat{i}_{k}}\big{)}.$ (3)
Let $N\in\mathbb{R}^{S\times A\times A}$ be a tensor that collects the proportions of each comparison in the dataset $\mathcal{D}$. In particular, $N_{k,i,j}\coloneqq\frac{1}{n}\sum_{\iota=1}^{n}\mathbbm{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}$ is the proportion of the number of times actions $i$ and $j$ have been compared under state $k$. A problem instance, denoted as $v$, is completely characterised by the tuple $(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$. In the following, we index instance-specific parameters with the instance $v$ to indicate their dependence on $v$; this will be omitted when the instance is clear from the context. In addition, we write $\mathbb{P}_{v}^{\pi}$ (resp. $\mathbb{E}_{v}^{\pi}$) to denote the probability measure (reps. the expectation) induced under policy $\pi$ and under the instance $v$. Other assumptions are as follows.
@@
In previous works (Zhu et al.,, 2023), the authors assume that the norms of the feature vectors $\phi(k,i)$ and parameter vector $\theta$ are separately bounded. This clearly implies that Assumption 2.1 is satisfied, but Assumption 2.1 is weaker as it is a bound on the rewards.
@@
We say an instance $v$ is inconsistent if it is not consistent. In the following, we will be only concerned with those instances that are consistent as the following result shows that it is impossible to design a policy that achieves vanishing simple regret for inconsistent instances.
@@
(Impossibility result) For any inconsistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$, there exists an instance $v^{\prime}=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta^{\prime})$ such that for all policies $\pi$

### v2

# On the Exponential Convergence for Offline RLHF with Pairwise Comparisons

@@ Abstract
We consider the problem of offline reinforcement learning from human feedback (RLHF) with pairwise comparisons proposed by Zhu et al. (2023), where the implicit reward is a linear function of an unknown parameter. Given an offline dataset, our objective consists in ascertaining the optimal action for each state, with the ultimate goal of minimizing the simple regret. We propose an algorithm, RL with Locally Optimal Weights or RL-LOW, which yields an exponential form of simple regret of $\exp(-\Omega(n/H))$ where $n$ is the number of data samples and $H$ denotes an instance-dependent hardness quantity that depends explicitly on the suboptimality gap of each action. Furthermore, we derive a first-of-its-kind instance-dependent lower bound in offline RLHF with pairwise comparisons. Interestingly, we observe that the lower and upper bounds on the simple regret match order-wise in the exponent, demonstrating order-wise optimality of our RL-LOW. In view of privacy considerations in practical applications, we also extend RL-LOW to the setting of $(\varepsilon,\delta)$-differential privacy and show, somewhat surprisingly, that the hardness parameter $H$ is unchanged in the asymptotic regime as $n$ tends to infinity; this underscores the inherent efficiency of RL-LOW in terms of preserving the privacy of the observed rewards. Given our focus on establishing instance-dependent bounds of exponential convergence, our research fills the research gap in existing studies that concentrate on establishing worst-case regrets of inverse polynomial convergence (e.g., $\widetilde{O}(\frac{1}{\sqrt{n}})$) for offline RLHF with pairwise comparisons.
@@ 1 Introduction
Reinforcement Learning (RL) (Sutton and Barto 2018) has been widely recognized for its capacity to facilitate agents in learning a sequence of optimal actions through iterative interactions with their environments. However, RL encounters significant hurdles in environments that are characterized by uncertainty or lacking explicit reward signals. To address these shortcomings, the concept of RL with human feedback (or RLHF) has emerged as a prominent paradigm. Preference-based RL (PbRL) (Christiano et al. 2017; Chen et al. 2022; Ibarz et al. 2018; Palan et al. 2019) has stood out as one of the most widely used frameworks for RLHF. In this regard, preference-based RL has achieved remarkable performances in practical applications, with particular importance lying in its ability to align large language models (LLMs) with human intent, thereby mitigating the output of toxic and dishonest information (Ouyang et al. 2022; Ziegler et al. 2019; Glaese et al. 2022; Bai et al. 2022; Liu et al. 2023), and improving the quality of applying to the specific tasks (Stiennon et al. 2020; Wu et al. 2021; Nakano et al. 2021).
@@
In this work, we tackle the problem of offline RLHF with pairwise comparisons, wherein the learning mechanism operates solely on pre-existing (or offline) data without dynamically engaging with the environment. Given the high cost associated with human interaction, offline RLHF has assumed particular importance in the context of incorporating human feedback. The significance of this offline framework has been justified by many previous prominent works (Shin et al. 2023; Ouyang et al. 2022; Zhu et al. 2023; Kim et al. 2023). For instance, within the learning process of InstructGPT (Ouyang et al. 2022) or the training procedure of Ahmadian et al. (2024), a pivotal procedure involves the training of a reward model utilizing pre-trained LLM feature vectors, coupled with the utilization of pre-collected human pairwise comparisons as the training dataset. Conceptually, this procedure can be construed as treating the current prompt context as a state within a certain Markov Decision Process (MDP), while the responses generated by the LLM serve as actions within this process. Empirical findings presented by Ouyang et al. (2022) demonstrate the efficacy of this offline framework in effectively aligning human intent with the outputs of LLMs.
@@
However, the literature concerning theoretical analyses within the domain of offline PbRL remains rather scant. Previous theoretical analyses (Zhu et al. 2023; Zhan et al. 2024) of offline PbRL predominantly focused on the worst-case (or minimax) regret, often resulting in the derivation of regret upper bounds for their algorithms of the form $\tilde{O}(n^{-1/2})$, where $n$ is the size of the offline dataset. In this work, we adopt a different approach that is centered on instance-dependent guarantees. In other words, we wish to derive performance guarantees that are functions of the specific problem instance, thus elucidating the role of fundamental hardness parameters. This yields complementary insights to the existing worst-case analyses. To this end, we design and analyze RL-LOW, a preference-based RL algorithm. Our analysis of the performance RL-LOW unveils an instance-dependent simple regret bound of $\exp(-\Omega(n/H))$, where $H$ is a hardness parameter. This reveals that the simple regret decays exponentially fast in the size of the dataset $n$ and the exponential rate of convergence has also been identified. Complementarily, by proving an instance-dependent lower bound, we show that any algorithm will suffer from a simple regret of at least $\exp(-O(n/H))$. Thus, the dependence of the problem on $H$ is fundamental and cannot be improved upon, thereby demonstrating the efficacy of RL-LOW and the tightness of our analyses.
@@ 1.1 Related Works
Preference-Based RL: From the empirical viewpoint, Christiano et al. (2017) initially demonstrated that RL systems can effectively address complex tasks like Atari games and simulated robot locomotion by learning from human preferences between trajectory segments. Later, numerous researchers started to employ human pairwise comparisons to enhance the performance of LLMs, e.g., aligning the LLMs’ behavior with human intent (Ziegler et al. 2019; Glaese et al. 2022; Bai et al. 2022; Liu et al. 2023), and enhancing the efficacy of application to specific tasks (Stiennon et al. 2020; Wu et al. 2021; Nakano et al. 2021).
@@
From the theoretical perspective, the existing literature remains sparse in offline RLHF with pairwise comparisons. Zhu et al. (2023) elucidated the failure of the maximum likelihood estimation (MLE) procedure in some scenarios. Motivated by this, they theoretically prove the (near) minimax optimality of the Pessimistic MLE approach with a high probability guarantee. In addition,  Zhan et al. (2024) introduced a novel paradigm for general reward functions, and they introduce $\varepsilon$-bracket approximations for reward models, accompanied by a rigorous theoretical analysis delineating sample complexity in terms of approximation error $\varepsilon$ and the high-probability parameter $\delta$. Recently, despite the significant contributions of Cen et al. (2024) and Liu et al. (2024) in advancing the integration of experimental findings and theoretical analysis in offline RLHF, the bounds they established are still characterized as worst-case bounds with inverse polynomial forms.
@@
We observe that the above theoretical investigations, while invaluable, are not instance-dependent and do not exhibit exponential convergence. Typically, the above minimax or worst-case guarantees yield upper bounds in the form of $\tilde{O}(n^{-1/2})$ and do not depend on any problem-specific factors (such as suboptimality gaps). Our research stands out as a pioneering attempt in offering an instance-dependent examination that yields exponential convergence for offline RLHF with pairwise comparisons, thereby bridging a critical gap in the existing literature.
@@
Label-Differential Privacy: In our study, we also consider the notion of label privacy, acknowledging that the labels in our offline dataset originate from users, thus highlighting the imperative to protect user privacy. Chaudhuri and Hsu (2011) were among the pioneers in exploring the concept of label privacy within the context of supervised learning for binary classification. Their foundational work posits that the sensitive information primarily resides in the labels, while considering the unlabeled attributes as non-sensitive. Later, the concept of label privacy has been investigated across various machine learning paradigms, including but not limited to PAC learning (Beimel et al. 2013) and deep learning frameworks (Ghazi et al. 2021). This broadened examination underscores the significance and relevance of label privacy considerations across diverse areas of machine learning research and applications.
@@
More recently, Chowdhury et al. (2024) investigated the use of label differential privacy to protect the privacy of human labelers in the process of estimating rewards from preference-based feedback. Chowdhury et al. (2024) derive an upper bound for their proposed algorithm on the estimation error. They show that it also decays as $O(n^{-1/2})$ and the implied constant here depends on $(\varepsilon,\delta)$, the parameters that define differential privacy. This bound only applies in the scenario of estimating the reward value and is not applicable if we want to understand how it depends on the simple regret of a specific instance. In our work, we consider the effect of $(\varepsilon,\delta)$-DP on the simple regret.
@@ 1.2 Our Contributions
    We extend RL-LOW to be applicable to the $(\varepsilon,\delta)$-differential privacy with labels by combining the Gaussian mechanism with the aforementioned locally optimal weights. Our analysis demonstrates that, for large datasets, this combination enables our algorithm to achieve differential privacy without weakening the bound on the simple regret, underscoring the superiority of the design and analysis of RL-LOW.
@@
    As a by-product of our analyses, we show that RL-LOW achieves a worst-case bound of the form $O(n^{-1/2})$ for the dependency of $n$. If we translate the high-probability upper bound in Zhu et al. (2023) to the same worst-case setting, we obtain a bound of the form $O(\sqrt{n^{-1}\log n})$. Thus, our work provides a noticeable (albeit small) improvement over the state-of-the-art theoretical result in (Zhu et al. 2023) in terms of the dependency on the sample size $n$.
@@ 2 Preliminaries and Problem Setup
Let $\mathcal{S}=\{1,\ldots,S\}$ denote the state space, and $\mathcal{A}=\{1,\ldots,A\}$ denote the action set. The $i$-th action of state $k$ is associated with the feature vector $\phi(k,i)\in\mathbb{R}^{d}$, and its associated (unknown) reward is
@@
where $\theta\in\mathbb{R}^{d}$ is an unknown parameter vector. The collection of all feature vectors is denoted as $\phi=\{\phi(k,i)\}_{k\in\mathcal{S},i\in\mathcal{A}}$. For all $k\in\mathcal{S}$, we denote the suboptimaliy gap of action $i\in\mathcal{A}$ as $\Delta_{k,i}=\max_{j\in\mathcal{A}}r_{k,j}-r_{k,i}$. Let $(a^{(0)},a^{(1)})\in\mathcal{A}^{2}$ be a pair of comparisons and let $s\in\mathcal{S}$ be a state. Then, we define a stochastic label $\sigma\in\{0,1\}$, following the Bradley–Terry–Luce (BTL) model as
@@
[…]
$R_{n}=\mathbb{E}_{k\sim\rho}\big[r_{k,i^{*}_{k}}-r_{k,\hat{i}_{k}}\big],$ (3)
where $\rho=(\rho_{1},\ldots,\rho_{S})$ is an unknown static distribution over states. Without loss of generality, we assume $\rho_{i}>0$ for $i\in\mathcal{S}$. We also consider a generalised version of the regret that is amenable to the MDP setting of RL in Section 5. by extending $\rho$ to be dynamic. Let $N\in\mathbb{R}^{S\times A\times A}$ be a tensor that collects the proportions of each comparison in the dataset $\mathcal{D}$, which satisefies $\lceil nN_{k,i,j}\rceil=\sum_{\iota=1}^{n}\mathbf{1}\{s_{\iota}=k,a_{\iota}^{(1)}=i,a_{\iota}^{(2)}=j\}$, where $N_{k,i,j}\in[0,1]$ is the proportion of the number of times actions $i$ and $j$ have been compared under state $k$. In this paper, we consider the concept of problem instance, denoted as $v$, which is characterised by $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$.
In the following, we index instance-specific parameters with the instance $v$ to indicate their dependence on $v$; this will be omitted when the instance is clear from the context. In addition, we write $\mathbb{P}_{v}^{\Pi}$ (resp. $\mathbb{E}_{v}^{\Pi}$) to denote the probability measure (reps. the expectation) induced under algorithm $\Pi$ and under the instance $v$. For notational brevity in the rest of the paper, we assume $nN_{k,i,j}\in\mathbb{N}$ is an integer.  [^2] For notational brevity, we assume that $nN_{k,i,j}\in\mathbb{N}$ is an integer. To be more precise, the sample count for $(k,i,j)$ should be written as $\lceil nN_{k,i,j}\rceil$.
@@
In previous works (Zhu et al. 2023), the authors assume that the norms of the feature vectors $\phi(k,i)$ and parameter vector $\theta$ are separately bounded. This clearly implies that Assumption 2.1 is satisfied, but Assumption 2.1 is weaker as it is a bound on the rewards.
@@
We say an instance $v$ is inconsistent if it is not consistent. In the following, we will be only concerned with those instances that are consistent as the following result shows that it is impossible to design a algorithm that achieves vanishing simple regret for inconsistent instances.
@@
(Impossibility Result) For any inconsistent instance $v=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta)$, there exists an instance $v^{\prime}=(\rho,\mathcal{S},\mathcal{A},\phi,N,\theta^{\prime})$ such that for all algorithms $\Pi$